Circular No. 6645

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) P. Colom, Observatoire de Paris (OP); D. Despois, Observatoire de

Bordeaux; B. Germain, OP; R. Moreno and G. Paubert, Institut de Radioastronomie Millimétrique (IRAM); N. Biver, D. Bockelée-Morvan, J. Crovisier, E. Gérard, E. Lellouch, and H. Rauer, OP; J. K. Davies, Joint Astronomy Centre, Hilo; and W. R. F. Dent, Royal Observatory, Edinburgh, report: “On Apr. 5 UT, we detected the blend of J(21–20) lines of CH_{3}OCHO (methyl formate) in C/1995 O1 at 227.562 GHz with the IRAM 30-m telescope. The total integrated line area in the main-beam brightness temperature scale is 0.22 ± 0.03 K km s^{−1}. A preliminary estimate of the CH_{3}OCHO production rate is 6 × 10^{27} molecules/s, which corresponds to an abundance relative to CH_{3}OH of ∼ 3 percent.”

R. W. Russell, D. K. Lynch, and A. L. Mazuk, The Aerospace Corporation, report the clear detection of features at 11.8 and 11.25 μm in the high signal-to-noise BASS spectra of comet C/1995 O1 in 1996 October and November, similar to features seen in the spectra of comet 1P/Halley. Spectral-synthesis modeling results in the identification of both features with emission by crystalline olivine grains, adding strength to their earlier detection (Bregman et al. 1987, A.Ap. 187, 616) and identification with crystalline olivine (Campins and Ryan 1989, Ap.J. 341, 1059) in 1P/Halley. In contrast to the earlier work, this model uses laboratory emissivity data on celestial dust-analogue samples as inputs. The presence of all of the expected structure in the region 8–13 μm strengthens the identification with a crystalline component of olivine-like composition in this comet.

GS 1843+00 K. Dennerl, Max-Planck-Institut für Extraterrestrische Physik, Garch-

ing; and J. Greiner, Astrophysical Institute, Potsdam, write: “This x-ray binary pulsar was very likely observed with the ROSAT High Resolution Imager (HRI) on Apr. 4.6608–4.6662 UT, when it had already declined back to the BATSE detectability threshold. A source was found at α = 18^{h}45^{m}36^{s}.9, δ = +0^{o}51′45′′ (equinox 2000.0; 90-percent confidence error radius 10′′), with an average countrate of 0.05 ± 0.01 HRI count/s. This source — the only object detected in this observation — is within the 2′ error radius of RXTE (cf. IAUC 6605). It was not present during an HRI observation on 1994 Apr. 8.5596–8.6348, which yields a 3σ upper limit of 8 × 10^{−4} HRI count/s. A direct identification with GS 1843+00 by detecting the 29.6-s pulsations is impossible due to limited photon statistics.”

1997 May 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green